OpenStudy (anonymous):

What is the sum of an 8-term geometric sequence if the first term is 10 and the last term is 781,250?

OpenStudy (amistre64):

the formula is: \[first\frac{1-r^n}{1-r}\]

OpenStudy (cwrw238):

first you need to find value of common ratio r r^7 = 782, 250 / 10 = 78250

OpenStudy (anonymous):

\[Sum = \frac{a(r^n - 1)}{(r-1)}\]

OpenStudy (anonymous):

\[r^7 = 78125\] Find r..

OpenStudy (cwrw238):

r = 5

OpenStudy (amistre64):

given that\[a_n=a_1*r^{(n-1)}\] \[\left(\frac{a_n}{a_1}\right)^{\frac{1}{n-1}}=r\] that was tough to type in lol

OpenStudy (anonymous):

i Got 976,560

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

\[Sum = \frac{10((5)^8 - 1)}{(5-1)}\]

OpenStudy (anonymous):

Well Done..

OpenStudy (anonymous):

You are right..

OpenStudy (anonymous):

thank you! all Of You!

OpenStudy (anonymous):

Welcome dear.